A box has a volume of 7500 cm3 and a surface area of 2600cm2. what are the length, width, and height of the box?
Found 2 solutions by fanism, xcentaur:
Answer by fanism(39)

you have 3 variables, but according to your post, there could only generate 2 equations. are u misssing something?

Answer by xcentaur(357)

let length be x cm,breadth be y cm,height be z cm


then volume=x*y*z=7500
then surface area=2(xy*yz*zx)=2600


We only get 2 equations,please check your question for any missing data or assumptions you are required to make,and post again
-xC